OpenStudy (anonymous):

Find the antiderivative: (x^2)^(1/3) + x(x)^(1/2)

OpenStudy (ipwnbunnies):

Easier if you simplify the integrand.

OpenStudy (campbell_st):

rewrite both terms using index laws \[(x^2)^{\frac{1}{3}} = x^{\frac{2}{3}}\] and \[x \times x^{\frac{1}{2}} = x^{\frac{3}{2}}\] now you can find the antiderivatives

OpenStudy (anonymous):

hmm ok thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I'm still a bit confused

OpenStudy (ipwnbunnies):

You can now rewrite it as \[\int\limits_{?}^{?} x^{2/3} + x^{3/2}\]

OpenStudy (anonymous):

Yes, so now

OpenStudy (anonymous):

I should have something like

OpenStudy (ipwnbunnies):

dx*

OpenStudy (anonymous):

(3/5)x^(5/3) + (2/5)x^(5/2)

OpenStudy (anonymous):

?

OpenStudy (ipwnbunnies):

There ya go.

OpenStudy (anonymous):

I took a break from doing this kind of problem and it's still difficult

OpenStudy (ipwnbunnies):

Don't forget to add C.

OpenStudy (anonymous):

Yes? that is it? ? :)))

OpenStudy (ipwnbunnies):

Yup. See, simplifying the integrand was the trick to the problem.

OpenStudy (anonymous):

coool

OpenStudy (anonymous):

thanks :)